Using the letters QRSTUVWXYZ, how many different arrangements could be made if X,Y, or Z must come first?
Thanks a ton for the help!
-Isaac

Answer by Edwin McCravy(20054)

Choose the first letter 3 ways,
Choose the 2nd letter 9 ways.
Choose the 3rd letter 8 ways.
Choose the 4th letter 7 ways.
Choose the 5th letter 6 ways.
Choose the 6th letter 5 ways.
Choose the 7th letter 4 ways.
Choose the 8th letter 3 ways.
Choose the 9th letter 2 ways.
Choose the 10th letter 1 ways.

3×9×8×7×6×5×4×3×2×1 = 3×9! = 1088640 ways.

Edwin